OpenStudy (omarbirjas):

WILL MEDAL write an equation for the translation of y=4/x that has the given asymptotes......x=0 and y=3

OpenStudy (omarbirjas):

@perl

OpenStudy (omarbirjas):

@Abhisar

OpenStudy (omarbirjas):

@dan815

OpenStudy (omarbirjas):

@e.mccormick @confluxepic

OpenStudy (omarbirjas):

@iambatman

OpenStudy (omarbirjas):

@Luigi0210

OpenStudy (omarbirjas):

@jagr2713 @Jaynator495

OpenStudy (omarbirjas):

OH FOR GODS SAKE SOMEBODY

OpenStudy (omarbirjas):

@Preetha

OpenStudy (omarbirjas):

@pooja195

OpenStudy (omarbirjas):

@vera_ewing

OpenStudy (canadiansteve99):

that's a picture of y=4/x. it's horizontal asymptote is y=0 and vertical one is x=0.since you want to move the vertical asymptote 7 units over to the right, your denominator will then be x-7... dnddsince you want the horizontal asymptote to move up 6 units, just add 6 to the y-values... so your equation will be: y = 1/(x-7) + 6

OpenStudy (canadiansteve99):

@JackofallTradez

OpenStudy (omarbirjas):

wait why is there a 1 in the numerator instead of a 4?

OpenStudy (omarbirjas):

how would you do it for x=-3 and y=-4

OpenStudy (canadiansteve99):

@JackofallTradez

OpenStudy (omarbirjas):

halp jack....

OpenStudy (omarbirjas):

he left ;-;

OpenStudy (omarbirjas):

@dan815

OpenStudy (omarbirjas):

write an equation for the translation of y=4/x that has the given asymptotes......x=0 and y=3 and then again for how would you do it for x=-3 and y=-4

OpenStudy (dan815):

making it y=4/(x-3) will give it a y=3 asymptote

OpenStudy (dan815):

sorry my bad

OpenStudy (dan815):

that will have a x=3 asymtote

OpenStudy (omarbirjas):

and for y we get y=4/x+0?

OpenStudy (dan815):